OpenStudy (ronaldo7):

7x+2 = x√5 Can anyone walk me through this thanks

OpenStudy (solomonzelman):

the right side is x times the square root of 5 ?

OpenStudy (solomonzelman):

or is it xTH root of 5?

OpenStudy (ronaldo7):

Yeah

OpenStudy (solomonzelman):

the first or the second?

OpenStudy (ronaldo7):

It is x times the square root of 5

OpenStudy (solomonzelman):

Example. \(\large\color{black}{ \displaystyle 8\cdot x+5 =\sqrt{19}\cdot x }\) \(\large\color{black}{ \displaystyle 8\cdot x-\sqrt{19}\cdot x+5 =0 }\) \(\large\color{black}{ \displaystyle (8-\sqrt{19})\cdot x+5 =0 }\) \(\large\color{black}{ \displaystyle (87-\sqrt{19})\cdot x=-5 }\) \(\large\color{black}{ \displaystyle x=-5/(8-\sqrt{19}) }\)

OpenStudy (ronaldo7):

im sorry im lost what did u do

OpenStudy (ronaldo7):

did u factor

OpenStudy (solomonzelman):

I subtracted the (√19)•x, then I factored the first two things out of x, then subtracted five from both sides, then divided by 8-(√19), {or by x's coefficient}, on both sides.

OpenStudy (ronaldo7):

is there any other way

OpenStudy (ronaldo7):

is there any other way

OpenStudy (solomonzelman):

what I did isn't that difficult (?)

OpenStudy (dayakar):

write all x terms leftside and constants right side